A project generates a cash flow of $499,900.00 per year (end of year cash flows). If the project can last 13.00 more years, what is its value TODAY of the remaining cash flows if the cost of capital is 9.00%?
Answer #1

Value of Project Today = PMT*(1-(1+r)-n)/r = 499,900*(1-(1+9%)-13/9% = 3,742,703.27

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well